# A Learning-Based Framework for Safe Human-Robot Collaboration with Multiple Backup Control Barrier Functions

### v1

@@ III-A Human Interface
Since we extract an intention estimation signal from the human driver, providing necessary feedback is essential for improving human-robot collaboration. Past work has suggested that systems which allow robots to estimate human intention, while their collaborating humans estimate robot intention, can improve overall performance [26]. Thus, we develop an interface based on the ROS rviz visualization tool as well as an Xbox controller featuring haptic motors. The rviz system displays a visualization of the robot pose, along with dials that indicate estimated system safety, as quantified by the magnitude of the saturated safety function $tanh(h(x))$. Furthermore, vibratory haptic feedback is provided to the user when the robot nears the boundary of the safe set as indicated by $h$. This feedback warns that the vehicle is in a potentially dangerous conditions, so we tune the the thresholds and strength of the haptic feedback as preferred by the driver.
@@ IV Implementation
Figure 3: Photo of our test-vehicle (a US Army DEVCOM GVR-Bot robot) operating in the NASA Jet Propulsion Laboratory Mars Yard. $(yellow\ inset)$ Intel Realsense D457 Depth Cameras are coupled with a VN-100 IMU, $(light-blue\ inset)$ custom compute payload.
Figure 4: Robot trajectory which used all three backup controllers under our learned switching law. Subfigure (a) shows the robot trajectory, where direction is indicated by the arrows and the choice of backup controller is indicated by color. Subfigures (b), (c), and (d) show the segments of the robot trajectory that used $\mathbf{k_{b1}},\mathbf{k_{b2}},\mathbf{k_{b3}}$ respectively. In (b), (c), and (d) we also show the flows of the respective backup controllers at several robot positions. Notice that in each of (b) (c) (d), the backup controller flow always escapes the obstacle. However, notice that $\mathbf{k_{b0}}$ may not provide safe evasion from the obstacle in the locations in subfigure (c) as it does for the locations in (b). This implies that our system chose the correct backup controller depending on several factors, like driver intent and robot position.
@@ IV-B Backup Controllers
We enforce safety in the presence of input bounds by implementing multiple BCBFs.
We leverage three backup controllers that yield qualitatively different behavior. Controller $\mathbf{k_{b0}}$ turns the robot away from the obstacle and drives away with maximum speed. Controller $\mathbf{k_{b1}}$ drives straight away as the obstacle is approached, without turning. Controller $\mathbf{k_{b2}}$ turns towards the obstacle and drives in reverse. It behaves similarly to $\mathbf{k_{b0}}$, however the robot has turned around. These are expressed by:
@@
Ultimately, the BCBFs enable the robot to maintain safe behavior. At the same time, the performance can be improved by switching between the backup policies based on learning. Safety must be preserved when switching between multiple backup controllers. Consider a switch from backup controller $\mathbf{k_{bi}}$ to $\mathbf{k_{bj}}$. One way to ensure safety during this crossover is to evaluate constraints (12) for the new controller $\mathbf{k_{bj}}$. We are guaranteed to be in the $T$-time reachable set of the first controller, $\mathbf{k_{bi}}$. Thus, validating that the BCBF-QP constraints hold for $\mathbf{k_{bj}}$ in practice requires that both backup controllers’ $T$-time reachable sets intersect. Furthermore, note that if backup controllers with $T$-time reachable sets are the only controllers switched between, the resulting QPs will always be feasible. This has the effect of, in practice, resulting in relatively smooth switching between controllers.
@@ V Results
Figure 5: Demonstration of system safety during switches between backup controllers. See that $h(x)\geq 0$ and the input constraints are satisfied (denoted by the gray dotted lines); therefore, our system maintains safety while better aligning with driver intention.
Figure 6: We compare our system on multiple environments with the use of a single backup controller, and also show the reproducibility of our test data. In figure (a), we attempt to drive the robot forwards to a point on the far right of the obstacle. In figure (b), we drive the robot and face the obstacle at various points. In figure (c), we attempt to reach the same goal as figure (a), but by driving the robot backwards. Notice our system provides more gentle corrections to driver controls and enables the driver to reach the goal position in every case.
By referring to Fig. 4, it is evident from the alternating colors that our learned reward function selected multiple backup controllers over a single robot trajectory. Our framework also selected appropriate backup controllers such that robot reachability near the obstacle was maximized. Furthermore, our system maintains the formal safety guarantees from the BCBF method as demonstrated in Fig. 5, where $h$ was observed to be positive under input limits. To demonstrate the reproducibility of these results, further tests were conducted and documented in Fig. 6 and the accompanying video shared in the caption of Fig 1. Here, we tested our system in many scenarios and compared produced trajectories to trajectories corresponding to a single backup controller. The use of our switching framework outperforms the use of a single backup controller in certain scenarios, and our system never inhibits the driver from achieving their goal, while some backup controllers do. For instance, sub-figure (b) of Fig 6 presents a case where $\mathbf{k_{b0}}$ inhibited the driver while our framework did not. The learned reward function correctly captures the driver’s intent, encompassing risk-neutral preferences such as emergency stopping and overtaking with minimal slowdown, as well as risk-averse behaviors like maintaining a sizeable safety distance.
@@ VI Conclusions and Future Work
Several next steps exist for this preliminary framework, like, developing an algorithm to formally compute safe reachability under multiple backup controllers. Furthermore, while we showed that our framework is useful on a teleoperated ground vehicle, we plan to implement our framework on other robots, such as quadrupeds, bipeds, and drones. Finally, various techniques can be employed to enhance the resilience of standard CBF constraint (6) against disturbance or model uncertainty, such as GP-based uncertainty in the CBF constraint [28], [29].
Acknowledgment. The authors would like to thank the SFP program at Caltech, SFP sponsors Kiyo and Eiko Tomiyasu, and DARPA for funding this project. Furthermore, the authors would like to thank Ryan K. Cosner and Maegan Tucker for discussions about safety with backup CBFs and learning, and Matthew Anderson for his insights and help in experimental setup and testing.

### v3

@@ III-C Human Interface
Since we extract an intention estimation signal from the human driver, providing feedback is essential for improving human-robot collaboration. Past work has suggested that systems which allow robots to estimate human intention, while their collaborating humans estimate robot intention, can improve overall performance [25]. Thus, we develop an interface based on the ROS rviz visualization tool as well as an Xbox controller featuring haptic motors. The rviz system displays a visualization of the robot pose, along with dials that indicate estimated system safety, as quantified by the magnitude of the saturated safety function $\tanh(h(x))$. Furthermore, vibratory haptic feedback is provided to the user when the robot nears the boundary of the safe set (indicated with $h$). We tune the the thresholds and strength of the haptic feedback as preferred by the driver.
Figure 3: Our test-vehicle (US Army DEVCOM GVR-Bot robot) in the NASA Jet Propulsion Laboratory Mars Yard. $(yellow\ inset)$ Intel Realsense D457 Depth Cameras are coupled with a VN-100 IMU, $(light-blue\ inset)$ custom compute payload.
@@ IV Implementation
The aforementioned framework was deployed in a simple obstacle avoidance task in a 2D environment. The robot position is measured from its center, so a buffer radius is given to the circular obstacle (a traffic cone) to account for half the length of the robot. The robot is driven to various locations around the environment to assess the accuracy of our framework to predict the desired backup strategy.
Figure 4: Robot trajectory which used all three backup controllers under our learned switching law. Subfigure (a) shows the robot trajectory, where direction is indicated by the arrows and the choice of backup controller is indicated by color. Subfigures (b), (c), and (d) show the segments of the robot trajectory that used $\mathbf{k_{b0}},\mathbf{k_{b1}},\mathbf{k_{b2}}$ respectively. In (b), (c), and (d) we also show the flows of the respective backup controllers at several robot positions. Notice that in each of (b) (c) (d), the backup controller flow always escapes the obstacle. However, notice that $\mathbf{k_{b0}}$ may not provide safe evasion from the obstacle in the locations in subfigure (c) as it does for the locations in (b). This implies that our system chose the correct backup controller depending on several factors, like driver intent and robot position.
@@ IV-B Backup Controllers
We enforce safety in the presence of input bounds by implementing multiple BCBFs. We leverage three backup controllers that yield qualitatively different behavior. Controller $\mathbf{k_{b0}}$ turns the robot away from the obstacle and drives forward. Controller $\mathbf{k_{b1}}$ drives straight away as the obstacle is approached, without turning. Controller $\mathbf{k_{b2}}$ turns towards the obstacle and drives in reverse. It behaves similarly to $\mathbf{k_{b0}}$, however the robot has turned around. These are expressed by:
@@
Ultimately, the BCBFs enable the robot to maintain safe behavior. At the same time, the performance can be improved by switching between the backup policies based on learning.
@@ V Results
By referring to Fig. 4, it is evident from the alternating colors that our learned reward function selected multiple backup controllers over a single robot trajectory. Our framework also selected appropriate backup controllers such that robot reachability near the obstacle was maximized. Furthermore, our system maintains the formal safety guarantees from the BCBF method as demonstrated in Fig. 5, where $h$ was observed to be positive under input limits. To demonstrate the reproducibility of these results, further tests were conducted and documented in the accompanying video, shared in the caption of Fig. 1. Here, we tested our system in many scenarios and compared our system’s produced trajectories to trajectories corresponding to a single backup controller. The use of our switching framework outperforms the use of a single backup controller in certain scenarios, and our system never inhibits the driver from achieving their goal, while some backup controllers do.
Figure 5: Demonstration of system safety during switches between backup controllers. See that $h(x)\geq 0$ and the input constraints are satisfied (denoted by the gray dashed lines); therefore, our system maintains safety while better aligning with driver intention.
@@ VI Conclusions and Future Work
Several next steps exist for this preliminary framework, like developing an algorithm to formally compute safe reachability under multiple backup controllers. Furthermore, while our framework was demonstrated on a teleoperated ground vehicle with stationary obstacles, we plan to deploy our framework on other robots, such as quadrupeds or drones, with moving obstacles. Finally, various techniques can be employed to enhance the resilience of standard CBF constraint (6) against disturbance or model uncertainty, such as GP-based uncertainty in the CBF constraint [27], [28].
Acknowledgment. The authors would like to thank the SFP program at Caltech, sponsors Kiyo and Eiko Tomiyasu, and DARPA for funding this project. The authors would also like to thank Ryan Cosner and Maegan Tucker for discussions about BCBFs and learning, and Matthew Anderson for his insights and help in experimental setup and testing.